OpenStudy (christos):

Solve f(x) = log4 (x+3) + log4(2-x)

OpenStudy (badhi):

this is a function, there is nothing to solve.

OpenStudy (mertsj):

BAdhi is correct but you could write it in another form.

OpenStudy (zehanz):

You can simplify it, however, using the rule \(log_{a}p+log_{a}q=log_{a}pq\).

OpenStudy (zehanz):

In ordinary words: the sum of the logarithms of two numbers is the logarithm of the product of these numbers. Please write down what your f becomes with this rule!

OpenStudy (anonymous):

\[\log _{a}m + \log _{a}n = \log _{a}(m \times n)\]

OpenStudy (anonymous):

\[f(x) = \log _{4} (x+3) + \log _{4}(2-x)\] \[f(x) = \log _{4}(x+3)(2-x)\]